# The Mean of Multi-Object Trajectories

### v1

@@ V Fréchet Mean via Gibbs Sampling
on $\mathbb{Y}$, where $\alpha$ is some positive constant. Since $e^{-\alpha(\cdot)}$ is bounded, and monotonically decreasing on $[0,\infty)$, the maxima of $\rho$ coincide with the minima of $g$ subject to the constraint $F(y)=1$ (this also holds if we replace $e^{-\alpha(\cdot)}$ by any bounded monotonically decreasing function). Hence, all samples from $\rho$ satisfy the constraint $F(y)=1$ and the most likely samples are the minima of $g$.
@@ V-A Gibbs Sampling for Mean Trajectory
[…]
@@
[…]
@@
Algorithm 6 Gibbs sampling for trajectory consensus with $G_{t}$ iterations. $\mathtt{{\small CatSampling}}$ stands for categorical sampling.
@@ VI Experimental Results
This section presents the first practical demonstration of Fréchet mean (FM) multi-object trajectory consensus, that is, the tractable computation of the geometric barycenter or “average” for multiple sets of trajectory valued elements. The results explore the behavior of FM consensus with OSPA(2) metric (defined in (18)), where $p=100$m, $c=100\sqrt{2}$m and $r=1$. Both of the proposed implementations are considered, namely greedy search as in Sections III-IV (with 5 and 100 iterations respectively for trajectory and multi-object trajectory consensus) as well as Gibbs sampling as in Section IV (with 100 and 10 iterations respectively for trajectory and multi-object trajectory consensus).
@@
In addition to deterministic estimates (wherein the Euclidean distance (ED) is used as the base distance), we also apply our technique to fuse multi-object trajectories with uncertainty. Presumably, the reported trajectories from each node are represented by Gaussian distributions, i.e., if a trajectory exists at time $k$, its state is a Gaussian distribution parameterized by a mean vector and a covariance matrix. We use Jeffrey’s divergence (JD) [^4] JD is a pseudo-metric since it does not satisfy the triangle inequality. as the base distance, and apply the technique proposed in [nielsen2009sided] to compute the respected centroids (also Gaussian distributions). Since computing the centroid Gaussian distributions is expensive, we only compute the Fréchet mean multi-object trajectory (with uncertainty) using the greedy search method.
A sample output from each of FM (Greedy), TC and DBSCAN-T is given in Figure 3. Further, we also illustrate the output from FM fusion with uncertainty via JD in Figure 4. It can be seen that outputs of the FM approaches correctly identify all trajectories with small delays in track initialization and termination. In contrast, the TC output shows some susceptibility to track switching and fragmentation, while DBSCAN-T additionally exhibits late initiation and termination. These observations are consistent with the fact that FM is a multi-scan or window-based formulation which processes batch data, whereas TC and DBSCAN-T are single-scan or recursive strategies which cannot change past estimates.
Figure 3: Sample outputs from FM with Euclidean distance and greedy search, TC and DBSCAN-T.
Figure 4: Sample output from FM fusion with uncertainty. The ellipses represent 95% confidence intervals of the Gaussian distributions. Dashed line ellipses are estimates from the nodes and solid ellipses are the fused results.
@@
The cardinality curve in Figure 5 indicates that FM consensus generally results in correct cardinality estimates with quick track initiation but some delay in track termination. This behavior reflects the output of the underlying tracking filters at each of the sensor nodes since they all exhibit the same characteristics which are reported as the consensus. The OSPA(2) curve in Figure 6 further indicates that the greedy search, using both ED and JD as base distances, and Gibbs sampling implementations have almost identical performance. Importantly, the FM results in the lowest OSPA(2) error of the three approaches since it implements window-based or multi-scan consensus scheme.
@@
Further analysis is presented in Table III which shows the final OSPA(2) error calculated over the entire scenario along with the total execution time. The final OSPA(2) errors for each of the node estimates are also shown in Table IV. The final OSPA(2) values confirm that FM consensus results in the lowest estimation error and significantly improves the raw sensor node outputs. Further, the greedy search implementation is marginally less accurate than Gibbs sampling but is orders of magnitude faster. The results suggest that the FM via greedy search produces significant improvements in consensus performance compared to existing approaches and with relatively efficient computation times.
@@
DBSCAN-T 39.88(5.50) 58.28(2.25)
TC 49.47(8.50) 385.87(12.95)
FM-ED (Greedy) 11.06(0.75) 37.06(3.60)
FM-JD  (Greedy) 11.45(1.08) 219.04(50.20)
FM-ED (Gibbs) 10.45(0.68) 1105.36(67.01)
@@ VII Conclusions
In this work, we developed a consensus notion for trajectories and multi-sets of trajectories using the Fréchet mean and OSPA-based metrics. We proposed methods for computing these means based on greedy search or Gibbs sampling. The greedy-search-based algorithms are numerically efficient and reach consensus much faster than state-of-the-art algorithms. In the absence of theoretical results on the convergence to the optimal solution, we provide empirical analysis by benchmarking the greedy-based solution against the computationally intensive Gibbs-sampling-based solutions that have well established convergence properties. Our experiments on distributed multi-object tracking have shown that the proposed consensus solutions significantly outperform state-of-the-art techniques. The greedy search algorithms outperform the current techniques both in terms of efficiency and accuracy by a significant margin, and are not far from the Gibbs-sampling-based solution in accuracy. Future research will focus on developing efficient algorithms for diverse tracking applications, including scenarios with unknown and non-common fields of view, where our consensus approach shows significant potential for adaptation and advancement.
@@ I Mathematical Proofs
I-A Proof of Proposition 1
@@
Let $u$ be a trajectory with the same domain as $\hat{v}$, i.e., $\mathcal{D}_{\hat{v}}=\mathcal{D}_{u}$, where $u(l)=\mu$, and $u(k)=\hat{v}(k)$ for each $k\in\mathcal{D}_{\hat{v}}-\left\{l\right\}$. Then, adding ${\displaystyle{\textstyle\sum_{n=1}^{N}}}\mathbf{1}_{\mathcal{\overline{D}}_{\!v^{(n)}\!}}^{(l)}c^{r}|\mathcal{D}_{u}\cup\mathcal{D}_{v^{(n)}}|^{-r}$ to both sides of the above inequality we have
[…]
@@
[…]

### v2

@@ V Fréchet Mean via Gibbs Sampling
on $\mathbb{Y}$, where $\alpha$ is some positive constant. Since $e^{-\alpha(\cdot)}$ is bounded, and monotonically decreasing on $0,\infty)$, the maxima of $\rho$ coincide with the minima of $g$ subject to the constraint $F(y)=1$ (this also holds if we replace $e^{-\alpha(\cdot)}$ by any bounded monotonically decreasing function). Hence, all samples from $\rho$ satisfy the constraint $F(y)=1$ and the most likely samples are the minima of $g$.
@@ V-A Gibbs Sampling for Mean Trajectory
$\gamma^{*}:=\gamma$; $\textrm{$\text{cost}^{*}$}:=$ $W^{(r)}(\gamma^{*})$;
@@
if $W^{(r)}(\gamma)<\textrm{$\text{cost}^{*}$}$ then $\gamma^{*}:=\gamma$; $\textrm{$\text{cost}^{*}$}:=W^{(r)}(\gamma)$;
@@
Algorithm 6 Gibbs sampling for trajectory consensus with $G_{t}$ iterations and categorical sampler (denoted as $\mathtt{{\small CatSampling}}$).
@@ VI Experimental Results
This section presents the first practical demonstration of Fréchet mean (FM) multi-object trajectory consensus, that is, the tractable computation of the geometric barycenter or “average” for multiple sets of trajectory valued elements. The results explore the behavior of FM consensus with the OSPA(2) metric (defined in (18)), where p=100p=100m and $c=100\sqrt{2}$m. We demonstrate the results for the first-order ($r=1$) and the second-order ($r=2$) Fréchet means, denoted as FM(1) and FM(2) respectively. Both of the proposed implementations are considered, namely greedy search as in Sections III-IV (with 5 and 100 iterations respectively for trajectory and multi-object trajectory consensus) as well as Gibbs sampling as in Section IV (with 100 and 10 iterations respectively for trajectory and multi-object trajectory consensus).
@@
Figure 4: Sample outputs from FM(1) with Euclidean distance and greedy search, TC and DBSCAN-T.
@@
Figure 5: Sample output from FM(1) fusion with uncertainty. The ellipses represent 95% confidence intervals of the Gaussian distributions. Dashed line ellipses are estimates from the nodes and solid ellipses are the fused results.
In addition to deterministic estimates (wherein the Euclidean distance (ED) is used as the base distance), we also apply our technique to fuse multi-object trajectories with uncertainty. Presumably, the reported trajectories from each node are represented by Gaussian distributions, i.e., if a trajectory exists at time $k$, its state is a Gaussian distribution parameterized by a mean vector and a covariance matrix. To illustrate the application of our technique in Fréchet-mean fusion of tracks with uncertainty, we use the techniques proposed in [nielsen2009sided] to compute the centroids of Gaussians w.r.t. Jeffrey’s divergence (JD) [^6] JD, which is a symmetrized KLD, is a pseudo-metric since it does not satisfy the triangle inequality. We use JD in our demonstration owing to the popularity of KLD for fusing distributions via the geometric mean.. Since computing the centroid Gaussian distributions is expensive, we only compute the Fréchet mean multi-object trajectory (with uncertainty) using the greedy search method.
A sample output from each of FM(1) (Greedy) via ED, TC and DBSCAN-T is given in Figure 4. Further, we also illustrate the output from FM(1) fusion with uncertainty via JD in Figure 5. It can be seen that outputs of the FM approaches correctly identify all trajectories. In contrast, the TC output shows some susceptibility to track switching and fragmentation, while DBSCAN-T additionally exhibits late initiation and termination. These observations are consistent with the fact that FM(1) is a multi-scan or window-based formulation which processes batch data, whereas TC and DBSCAN-T are single-scan or recursive strategies which cannot change past estimates.
@@
The cardinality curves in Figure 6 indicate that FM consensus generally results in correct cardinality estimates with quick track initiation but some delay in track termination. This behavior reflects the output of the underlying tracking filters at each of the sensor nodes since they all exhibit the same characteristics which are reported as the consensus. The OSPA(2) curves in Figure 7 further indicate that the greedy search (for both FM(1) and FM(2)), using either ED and JD as base distances, and Gibbs sampling implementations have almost identical performance. Importantly, the FM results in the lowest OSPA(2) error of the three approaches since it implements window-based or multi-scan consensus scheme.
@@
Further analysis is presented in Table III which shows the final OSPA(2) error calculated over the entire scenario along with the total execution time. The final OSPA(2) errors for each of the node estimates are also shown in Table IV. The final OSPA(2) values confirm that FM consensus results in the lowest estimation error and significantly improves the raw sensor node outputs. Further, the greedy search implementation is marginally less accurate than Gibbs sampling but is orders of magnitude faster. Among the FM consensus algorithms, the second-order FM techniques exhibit slightly worse performance than their first-order counterparts, while demonstrating a noticeable reduction in run-time. The better performance of the order-1 FM could be attributed to the resilience of the median against noise. The more efficient order-2 FM (with the Euclidean base distance) is due to the simpler computation of the fused trajectory state at each time step by averaging the corresponding data points, whereas computing the median requires an iterative process. Overall, the results suggest that the FM via greedy search produces significant improvements in consensus performance compared to existing approaches and with relatively efficient computation times.
@@
DBSCAN-T 38.05(5.38) 39.19(0.93)
TC 49.47(8.50) 397.36(7.24)
FM(1)-ED (Greedy) 11.05(0.80) 27.57(2.29)
FM(2)-ED (Greedy) 12.05(1.05) 21.55(1.66)
FM(1)-ED (Gibbs) 10.48(0.81) 1099.52(76.84)
FM(2)-ED (Gibbs) 11.21(0.81) 1073.81(61.44)
FM(1)-JD  (Greedy) 11.41(1.06) 154.47(17.59)
@@ VII Conclusions
In this work, we introduced the notion of average for trajectories and multi-sets of trajectories using the Fréchet mean and OSPA-based metrics. We proposed methods for computing these means based on greedy search or Gibbs sampling. The greedy-search-based algorithms are numerically efficient and reach consensus much faster than state-of-the-art algorithms. In the absence of theoretical results on the convergence to the optimal solution, we provide empirical analysis by benchmarking the greedy-based solution against the computationally intensive Gibbs-sampling-based solutions that have well established convergence properties. Our experiments on distributed multi-object tracking have shown that the proposed consensus solutions significantly outperform state-of-the-art techniques. The greedy search algorithms outperform the current techniques both in terms of efficiency and accuracy by a significant margin, and are not far from the Gibbs-sampling-based solution in accuracy. Future research will focus on developing efficient algorithms for diverse tracking applications, including scenarios with unknown and non-common fields of view, where our consensus approach shows significant potential for adaptation and advancement.
@@ I Mathematical Proofs
I-A OSPA Trajectory Metric Proof
Using the abbreviation $\mathcal{D}_{\!uv}=\mathcal{D}_{\!u}\cup\mathcal{D}_{\!v}$, we rewrite the OSPA distance (2) between any two trajectories $u,v\in\mathbb{T}$ as
$d_{\mathbb{T}}^{(c,r)}(u,\!v)\!=\!\left[\frac{1}{|\mathcal{D}_{\!uv}|}\sum_{k\in\mathcal{D}_{\!u}\cup\mathcal{D}_{\!v}}[\tilde{d}_{\mathbb{X}}^{(c)}\!(u_{k},v_{k})]^{r}\right]^{\frac{1}{r}},$
where $u_{k}=\{u(k)\}$ if $k\in\mathcal{D}_{\!u}$, $u_{k}=\emptyset$ if $k\notin\mathcal{D}_{\!u}$,
$\tilde{d}_{\mathbb{X}}^{(c)}(x,y)=\begin{cases}d_{\mathbb{X}}^{(c)}(x^{(1)},y^{(1)}),&|x|=|y|=1\\ 0,&|x|=|y|=0\\ c,&\textrm{otherwise}\end{cases},$
and $x\!=\!\{x^{(1)}\!,...,x^{(n)}\!\}$. The proof is based on the following lemma.
Lemma A1. For any trajectories $u$, $v$, $t\in\mathbb{T}$, $c>0$ and $r\in[1,\infty]$, ,
$\left[\frac{1}{|\mathcal{D}_{\!ut}|}\!\sum_{k\in\mathcal{D}_{\!ut}}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})]^{r}\right]^{\frac{1}{r}}\!\!\!\leq\!\!\left[\frac{1}{|\mathcal{D}_{\!uv}|}\!\sum_{k\in\mathcal{D}_{\!uv}}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})]^{r}\!\right]^{\frac{1}{r}}\!\!.$
Proof. For $r=\infty$, the inequality holds because
$\lim_{r\rightarrow\infty}\left[\frac{1}{n}\sum_{i=1}^{n}[\alpha_{i}]^{r}\right]^{\frac{1}{r}}=\max_{1\leq i\leq n}\alpha_{i},$ (A1)
for $\alpha_{i}\geq 0$ and $\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})=0$ for $k\notin\mathcal{D}_{\!uv}$.
For $r\in[1,\infty)$, let $I=\mathcal{D}_{\!uv}\cap\mathcal{D}_{\!ut}$ and $V=\mathcal{D}_{\!uv}\setminus\mathcal{D}_{\!ut}$, which means $\mathcal{D}_{\!uv}=I\uplus V$. Noting that $|\mathcal{D}_{\!uv}|=|I|+|V|$ and $|I|=|\mathcal{D}_{\!uv}\cap\mathcal{D}_{\!ut}|\leq|\mathcal{D}_{\!ut}|$ we have,
$|\mathcal{D}_{\!uv}|\thinspace|I|\leq\thinspace$ $|\mathcal{D}_{\!ut}|(|I|+|V|)$
$\Leftrightarrow(|\mathcal{D}_{\!uv}|-|\mathcal{D}_{\!ut}|)|I|c^{r}\leq\thinspace$ $|\mathcal{D}_{\!ut}||V|c^{r}$
$\Leftrightarrow(|\mathcal{D}_{\!uv}|-|\mathcal{D}_{\!ut}|)\sum_{k\in I}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})]^{r}\leq\thinspace$ $|\mathcal{D}_{\!ut}||V|c^{r}$
$\Leftrightarrow\!|\mathcal{D}_{\!uv}|\sum_{k\in I}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})]^{r}\!$ $\!\leq\!|\mathcal{D}_{\!ut}|\!\left[\sum_{k\in I}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})]^{r}\!+|V|c^{r}\right]$
$\!\!\!\!\!\!\Leftrightarrow\!\frac{1}{|\mathcal{D}_{\!ut}|}\!\sum_{k\in I}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})]^{r}\!$ $\!\leq\!\frac{1}{|\mathcal{D}_{\!uv}|}\!\!\left[\sum_{k\in I}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})]^{r}\!\!+|V|c^{r}\!\right]\!\!.$
Since $\mathcal{D}_{\!ut}=I\uplus(\mathcal{D}_{\!ut}\setminus\mathcal{D}_{\!uv})$ and $\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})=0$ for $k\notin\mathcal{D}_{\!uv}$,
$\sum_{k\in\mathcal{D}_{\!ut}}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})]^{r}=$ $\sum_{k\in I}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})]^{r}.$
In addition, since $\mathcal{D}_{\!uv}=I\uplus V$ and $\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})]^{r}=c^{r}$ for $k\in V$ (i.e., $k\in\mathcal{D}_{\!v}$, $k\notin\mathcal{D}_{\!u}$, $k\notin\mathcal{D}_{t}$ ),
$\sum_{k\in\mathcal{D}_{\!uv}}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})]^{r}=$ $\sum_{k\in I}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})]^{r}+|V|c^{r}.$
Hence,
$\frac{1}{|\mathcal{D}_{\!ut}|}\sum_{k\in\mathcal{D}_{\!ut}}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})]^{r}\leq\frac{1}{|\mathcal{D}_{\!uv}|}\sum_{k\in\mathcal{D}_{\!uv}}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})]^{r}.$
Raising both sides to $1/r$ yields the desired result. $\blacksquare$
By definition, (2) trivially satisfies the metric properties except for the triangle inequality. Thus, for any trajectories $u$, $v$, $t\in\mathbb{T}$, we need to prove
$d_{\mathbb{T}}^{(c,r)}(u,v)\leq d_{\mathbb{T}}^{(c,r)}(v,t)+d_{\mathbb{T}}^{(c,r)}(u,t).$ (A2)
Consider $r\in[1,\infty)$. Since $\tilde{d}_{\mathbb{X}}^{(c)}$ is a metric,
$\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})\leq\tilde{d}_{\mathbb{X}}^{(c)}(v_{k},t_{k})+\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},t_{k}),$
and hence,
$\left[\frac{1}{|\mathcal{D}_{\!uv}|}\sum_{k\in\mathcal{D}_{\!uv}}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},v_{k})]^{r}\right]^{\frac{1}{r}}$
$\leq$ $\left[\frac{1}{|\mathcal{D}_{\!uv}|}\sum_{k\in\mathcal{D}_{\!uv}}[\tilde{d}_{\mathbb{X}}^{(c)}(v_{k},t_{k})+\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},t_{k})]^{r}\right]^{\frac{1}{r}}.$
Using the definition of $d_{\mathbb{T}}^{(c,r)}$ and applying Minkowski’s inequality to the right-hand side, it follows that
$d_{\mathbb{T}}^{(c,r)}(u,v)$ $\leq\left[\frac{1}{|\mathcal{D}_{\!uv}|}\sum_{k\in\mathcal{D}_{\!uv}}[\tilde{d}_{\mathbb{X}}^{(c)}(v_{k},t_{k})]^{r}\right]^{\frac{1}{r}}$
$+\left[\frac{1}{|\mathcal{D}_{\!uv}|}\sum_{k\in\mathcal{D}_{\!uv}}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},t_{k})]^{r}\right]^{\frac{1}{r}}.$
Applying Lemma A1 yields
$\left[\frac{1}{|\mathcal{D}_{\!uv}|}\sum_{k\in\mathcal{D}_{\!uv}}[\tilde{d}_{\mathbb{X}}^{(c)}(v_{k},t_{k})]^{r}\right]^{\frac{1}{r}}\!\!\!$ $\leq\!\!\left[\frac{1}{|\mathcal{D}_{\!vt}|}\sum_{k\in\mathcal{D}_{\!vt}}[\tilde{d}_{\mathbb{X}}^{(c)}(v_{k},t_{k})]^{r}\right]^{\frac{1}{r}}\!\!\!,$
$\left[\frac{1}{|\mathcal{D}_{\!uv}|}\sum_{k\in\mathcal{D}_{\!uv}}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},t_{k})]^{r}\right]^{\frac{1}{r}}\!\!\!$ $\leq\!\!\left[\frac{1}{|\mathcal{D}_{\!ut}|}\sum_{k\in\mathcal{D}_{\!ut}}[\tilde{d}_{\mathbb{X}}^{(c)}(u_{k},t_{k})]^{r}\right]^{\frac{1}{r}}\!\!\!.$
Thus, (A2) holds by the definition of $d_{\mathbb{T}}^{(c,r)}$ and the above inequality.
For $r=\infty$, it follows from (A1) that $\lim_{r\rightarrow\infty}d_{\mathbb{T}}^{(c,r)}(u,v)=d_{\mathbb{T}}^{(c,\infty)}(u,v)$. Thus, (A2) also holds using the same argument as per the $r\in[1,\infty)$ case.
I-B Proof of Proposition 1
@@
${\displaystyle\sum_{n:\mathcal{D}_{\!v^{(n)}\!}\ni l}}\frac{[d_{\mathbb{X}}^{(c)\!}(\mu,v^{(n)\!}(l))]^{r}}{|\mathcal{D}_{\hat{v}}\cup\mathcal{D}_{v^{(n)}}|}$ $\!\!\!<\!\!$ ${\displaystyle\sum_{n:\mathcal{D}_{\!v^{(n)}\!}\ni l}}\frac{[d_{\mathbb{X}}^{(c)\!}(\hat{v}(l),v^{(n)\!}(l))]^{r}}{|\mathcal{D}_{\hat{v}}\cup\mathcal{D}_{v^{(n)}}|}.$
@@
$\Psi_{l}^{(r)}(u)\!\!$ $\!\!\!\!=\!\!\!\!$ $\!\!{\displaystyle\sum_{n=1}^{N}}\frac{\mathbf{1}_{\mathcal{D}_{\!v^{(n)}\!}}^{(l)}[d_{\mathbb{X}}^{(c)\!}(\mu,v^{(n)\!}(l))]^{r}+\mathbf{1}_{\mathcal{\overline{D}}_{\!v^{(n)}\!}}^{(l)}c^{r}}{|\mathcal{D}_{u}\cup\mathcal{D}_{v^{(n)}}|}$
@@
$\!\!\!\!\!\!V^{(r)}(u)$ $\!\!\!=\!\!\!$ ${\displaystyle\sum_{k\in\mathbb{K}}}\mathbf{1}_{\mathcal{D}_{u}}^{(k)}\Psi_{k}^{(r)}(u)+\mathbf{1}_{\mathcal{\overline{D}}_{u}}^{(k)}\bar{\Psi}_{k}^{(r)}(u)$